OpenStudy (anonymous):

What is the transformation that occurs to the equation y = 2x if it changes to y = -2x − 2?

OpenStudy (lgbasallote):

can you write it more clearly?

OpenStudy (anonymous):

2^x if it changes to y=-2^x-2

OpenStudy (anonymous):

like that?

OpenStudy (anonymous):

this may help \[a^{b-c}=a^ba^{-c}\]

OpenStudy (anonymous):

\[-(2^{x-2}=-(2^x2^{-2})=-\frac{2^x}{2^2}\]

OpenStudy (anonymous):

oh ok that what you meant xD I was like.. uh...... hahah.. ok anyway

OpenStudy (anonymous):

or wait is it this \[2^x-2\]

OpenStudy (anonymous):

the bottom one

OpenStudy (anonymous):

ahh ok this makes more sense. Alright so find the y intercept by making x=0, what do you get

OpenStudy (anonymous):

i got the answer now, thanks ^_^